Find the expected value of the random variable X and its variance having the following density function:


f(x) = 5 (1-x4) 0<x<2


= 0 elsewhere


Full details math and explain.

1
Expert's answer
2022-03-23T19:07:33-0400

Expected value

"E(X)=5∫_0^2x(1-x^4 )dx=5∫_0^2x-x^5 dx=5[\frac{x^2}{2}-\frac{x^6}{6}]_0^2=5[(\frac{2^2}{2}-\frac{2^6}{6})-0]=-\frac{130}{3}"


Variance

"Var(X)=E(X^2)-(E(X))^2"

"E(X^2)=5∫_0^2x^2 (1-x^4 )dx=5∫_0^2x^2-x^6 dx=5[\frac{x^3}{3}-\frac{x^7}{7}]_0^2=5[(\frac{2^3}{3}-\frac{2^7}{7})-0]=-\frac{1640}{21}"

"Var(X)=-\frac{1640}{21}-(-\frac{130}{3})^2=-\frac{123220}{63}=-1955.8730"